Circular No. 8417

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004eu IN MCG +07-5-39 T. Boles, Coddenham, England, reports the discovery of an appar-

ent supernova (mag 17.3) on unfiltered CCD images taken on Oct. 6.203 and 6.890 UT with a 0.35-m reflector. The new object is located at α = 2^{h}20^{m}34^{s}.86, δ = +41^{o}34′18′′.5, which is ≈ 2′′.1 west and 7′′.9 south of the center of MCG +07-5-39. SN 2004eu is not present on Boles’ images from 2004 Sept. 15 and 2003 Dec. 28 (limiting mag 19.5) or on Digitized Sky Survey red (1987 Oct. 18, limiting mag 21.0) and blue plates (1989 Sept. 6, limiting mag 21.5).

SUPERNOVA 2004ev IN ESO 459-G13 Further to IAUC 8405, T. Puckett and J. Newton report the discovery

of an apparent supernova (mag 18.1) on an unfiltered CCD frame taken with the 0.35-m automated supernova patrol telescope on Oct. 6.03 UT, confirmed by Puckett on a frame taken on Oct. 7.03 (at mag 18.1) with a 0.60-m reflector. The new object is located at α = 19^{h}20^{m}31^{s}.78, δ = −28^{o}55′14′′.9 (equinox 2000.0), which is 21′′.4 east and 3′′.8 south of the center of ESO 459-G13. The candidate is not present on images taken by Puckett on Sept. 19 (limiting mag ∼ 20.0) or on Palomar Sky Survey plates taken on 1992 May 29 (limiting mag ∼ 21.0) and 1974 July 23 (limiting mag ∼ 20.0).

V1187 SCORPII R. J. Rudy, D. K. Lynch, and R. W. Russell, The Aerospace Corpo-

ration; and S. J. Bus, University of Hawaii, report infrared spectroscopy of V1187 Sco between 0.8 and 5 μm using the Infrared Telescope Facility (+ SpeX). The object entered its coronal phase around the middle of September, with the appearance of the [S VIII] and [S IX] lines on Sept. 22. Further observations on Sept. 30 revealed more and strengthening coronal lines. Currently present are [S VIII] at 0.99 μm and [S IX] at 1.25 μm, [Si VI] and [Si VII] at 1.96 and 2.48 μm, and weak [Ca VIII] at 2.32 μm. Despite the existence of coronal features and the emission lines of He II, fluorescently excited lines of O I persist. Brα is strong, and there are weak, unidentified features at 3.82 and 3.91 μm with broad flat-topped or castellated profiles similar to the hydrogen lines. There is no evidence that dust formation has occurred.

2004 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green